A 4 step process was established:
*Remove the wreckage (done June 11th, 2012);
*Secure the optics and telescope (done June 13/14, 2012);
*Close up the open dome room area against precip (done 3 July 2012);
*Install a replacement dome (20 Oct 2012) - a winter cap, but not yet operable.
Summer 2014: During 2013, a series of problems resulted in our contractor being unable to deliver on contractually agreed items, leaving the new dome non-functional. Because of the short summer access season, we were not able to resolve these matters last year, but geared up to address same during summer 2014. Primary upgrades needed include the dome slit lift mechanisms, and the dome rotation mechanisms. As of Memorial Day 2014, CDOT had completed plowing to the summit in time for Memorial Day traffic, despite the heavier-than-average snow season. During June, we have re-activated dome rotation and begun design and measurements for the slit door motor assembly. During July, slit lift machinery was constructed and installed during August, leading to a test run in late summer. During that process, we quantified the problem with dome rotation: the dome is not level. The following measurements were obtained during Sept 2014, for several of the 20 roller wheel assemblies, including support wall to gear track vertical distance, and support wall to gear track horizontal offset (inboard or outboard):
|Cardinal direciton||Support wheel set (marked)||Support wall to gear track vertical ht||Support wall face to gear track horizontal offset, inboard or outboard|
|NW||5||2.0 inches||0.5 inch out|
|WNW||7||2.1 in||0 (flush)|
|SW||11||2.2 in||0.5 inch inboard|
|S||14||1.6 in||0 (flush)|
|S||15||1.6 in||0 (flush)|
|ENE (near stairs)||20||2.0 in||0.5 in out|
Outside views confirm that the dome/skirt is tightest to the cylinder wall toward the SE direction, essentially tilted. Whether the excessive friction to rotation is due to wheel or wheel cradle interference, remains to be determined, next summer.
Summer 2013: Goals: obtain dome functionality & restore telescope. 11 July 2013: A crucial step completed, wiring repair enabled telescope to move and interface with Sky software control once again. Telescope rotation also revealed flaws in the cooling system (minor spills). Also during July 2013: Contractor discovers problems with out-of-round shape of dome cylinder wall, causing binding with dome skirt plates. This will require bumping out plate supports for a portion of the azimuth sweep. This lead to a series of delays and re-design, when combined with contractor schedule issues, meant that we only got the dome skirt plate installed and adjusted, plus a dome rotation motor assembly installed but not tested before autumn road closures. As of this writing, the following items will need to be attacked during summer 2014: (a) dome support wall out of round by 1 inch or more, enough to risk breaking wheels unless guide rollers are installed; (b) dome rotation motor control untested - to be done; (c) slit door presently bolted shut - need to design and install lift motor with manual override; (d) address superficial damage caused by initial dome failure plus contractor errors; (e) ideally, schedule a telescope use class for late summer 2014.
- - - - - HISTORY OF THE EVENTS, 2011 - 2012 - - - - -
NOV.2011: 11th - outdoor webcam** view shifts from horizontal to staring at the ground, possibly associated with major windstorm that opened the dome and snapped old growth trees at Doolittle Ranch below Echo Lake, as reported by owner Bob Mischke. ** http://netserv.du.edu/video/webpagedump/MTEVANS2.jpg A resident of Silver Plume reports measuring a 132 mph wind gust. Widespread blowdown of trees NW of mile 4 on Mt.Evans highway (noted later on by Aaron Reid).
- - - - -
DEC.2011/JAN.2012 - high wind speeds noted for 12/31 and especially 1/17-18 when sustained hourly averages above 90mph were estimated lasting for nearly a whole day.
- - - - -
FEB.2012 : 10th - initial contact from Adam Jones, skilled mountain climber, who wished to practice hike at/around Mt.Evans observatory in prep for his upcoming Mt.Everest summit attempt: see http://www.mountwashington.com/everest/ & http://www.mountwashington.com/bierstadt/
11th - another hiker, then atop Mt.Bierstadt, posts image of Mt.Evans summit, which on close inspection shows dome damage and exposed telescope.
- - - - -
MARCH 2012: 6th - webcam view shows shadow of dome looking peculiar as sun angle changes - openings in roof allow sunlight to show through...
15th - Adam Jones reaches summit on foot, provides first hand account and photos confirming suspected damage - see: http://mysite.du.edu/~rstencel/MtEvans/RepairSummary.pdf
23rd, 30th - subsequent visits by Adam Jones stabilize and secure the telescope, debris collected and stashed, other equipment moved to safety.
- - - - -
APRIL 2012 : 6th - final site visit by Adam Jones prior to his departure for Nepal and Mt. Everest.
- - - - -
MAY 2012: 1st - CDOT reports road cleared of snow to the summit.
16th - first access by vehicle and visual inspection of damage. Yes, as bad as it seemed in Adam's photos. New concern about damage to the worm and drive gears due to dome weight dropped on west arm of the mount.
17th - purchase request drafted for mirror cleaning and recoating.
18th - meeting with contractors to discuss next steps.
- - - - -
JUNE 2012 : 1st - demolition bid selection meeting, further discussion of dome re-design, emergence of financial issues.
6th - JP Meyer Trucking selected for dome demo. Site visit today to establish safety perimeter and clean up slit door wreckage from downhill areas to the NE. 11th - crane-lift dome wreckage removal accomplished.
13th - secondary mirror assemblies removed.
14th - primary mirror assemblies removed.
Week of 18th - temporary roofing installation started week of 25th - progress on temporary roofing, but high winds slow the pace.
- - - - -
JULY 2012: 3rd - temporary roof enclosure completed - ...see http://mysite.du.edu/~rstencel/MtEvans/Pix/PICT0151.JPG
15th - new dome design concept finalized.
26th - useful discussions with Tom Melsheimer, telescope automation genius, points out the need to: 1. address thermal expansion issues with new dome -> lots of stainless steel pop rivets in new skin; 2. motors need to be spring loaded to ride the variations in the gear rack radius ; 3. motors need an accel/decel ramp to avoid shock of too much inertia & he helped articulate the difference between MWO and other high end facilities - we are more experimental rather than specific project oriented. Next step: what are the important experiments?
27th - initial steps of fabrication of new dome.
- - - - -
AUG.2012: Dome fabrication proceeding. ...see * http://mysite.du.edu/~rstencel/MtEvans/Pix/061b.JPG
- - - - -
SEPT.2012: 4th - first fit check of slit door on dome rib structure - looks good.
14th - new parts needed for adding exterior skin of 1/8th inch sheet steel also first snowfall, summit - race against time...
- - - - -
2012 OCT.3: Dome fabrication nearing completion - almost ready to paint and ship ...see * http://mysite.du.edu/~rstencel/MtEvans/Pix/061c.JPG - Road has remained largely high and dry, the drought continues...
OCT.20 - somehow the road has remained passable past the usual early Oct snow closures of prior years, and as a result, NEW DOME DELIVERED! See: http://mysite.du.edu/~rstencel/MtEvans/Pix/061d.JPG
Site visits continued during early Nov - unusually dry road. Hardly the end of the story - still to be done: weatherstripping, motor controls for dome rotation and opening, dealing with possible telescope damage from the original event. Much of this may have to wait for summer 2013 access season. Thanks for your continued interest. Check back for updates.
WINTER 2012/13: Dormancy. As seen on our weather station site, http://gpsmet.noaa.gov/cgi-bin/gnuplots/rti_plots.cgi?SITES=COME the cold arrived on Dec.9th with a drop to -20F (-30C), followed by a month barely above 20F. The next substantial cold wave started Jan.11th, staying close to -20F for several days in a row...
SPRING 2013: snowy . Finally in late
May, the snows have stopped and the weather has warmed. From campus,
we can see the CDOT crews clearing snow on highway 5, and they might
summit by Memorial Day (5/24) afterall. Our planning for repairs
centers on getting the dome functional, then repairing the telescope.
One complication involves our contractor being heavily committed to
other projects. Stay tuned.
- - - - -
JUNE 2013: First 2 weeks featured removal of the snow and ice on the observatory access road, to enable supplies to arrive at the building.
Wed June 19: first sets of scaffolding arrive and are set up, but mismatched parts slow this process.
Mon June 24: lift crane on site to recenter the dome and level the wheels. Result is that the 22,000 lb. dome now can be turned. Several benign weather days missed due to contractor scheduling issues.
- - - - -
JULY 2013: welding and motorization next steps, then telescope rehab... see top of page for continued story.
Photos available on request.
In review, 2012: Narrowly, we had to remove the hazardous material that could have fallen on public nearby, and move to protect the twin telescope that was fully exposed to high altitude weather. More broadly, with University support, I feel obliged to return the facility to operational condition whether or not we can find my eventual successor to direct it, and because it is worth sustaining as one of the very few high altitude astronomy sites in North America. More broadly still, whilst I can and should be helping to resolve the mysteries of the binary star epsilon Aurigae**, the charge of my endowed professorship also includes the astronomy infrastructure and observatories. Though I've enjoyed my time in this role, it seems providing opportunities for students and promoting what educational research opportunities might exist in this "industrial astronomy" contnues to have merit, if we believe that there is a future for ground based astronomy. Thus, full repair and return to program seems like the proper course at this juncture.
see also bookend: http://adsabs.harvard.edu/abs/1999JAVSO..27...61S .